OpenStudy (anonymous):

Find the lateral area, total area, and volume for the cylinder with the given measurement. r = 6, h = 20

OpenStudy (anonymous):

Lateral area = 2pi (6x 20) Surface area = 2 pi (6 x 6) + 2 pi (6 X 20) Volumen = pi (6 x 6) x 20

OpenStudy (anonymous):

what is your answer for each of these if you calculate them ?

OpenStudy (anonymous):

Lateral = 2pi x 120 Surface area = 2pi x 36 + 2 pi x 120 Volume = pi x 36 x 20 Does this help break down the equation for you?

OpenStudy (anonymous):

a lil